OpenStudy (anonymous):

If sin(θ) = -2/√29, find sin^4(θ).

OpenStudy (anonymous):

(-2/sqrt(29))^4 = (-2)^4 / (sqrt(29))^4 = 16 / 29^2 = 0.0190